OpenStudy (anonymous):

What is the value of n in the equation 8n + 6 = 10? Please enter your answer as a fraction.

OpenStudy (anonymous):

8n + 6 = 10 Subtract 6 from both sides 8n = 4 Are you with me so far?

OpenStudy (anonymous):

Yes, it would be 2n = 4?

OpenStudy (anonymous):

You will then divide each side by 8 8n/8 = 4/8 The 8s cancel each other out n = 4/8 n=1/2

OpenStudy (anonymous):

I would not be 2n =4 The 8 is connected to the variable. You don't touch that until the end.

OpenStudy (anonymous):

Ah, thanks. I understand it now. Medal for you! :-)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

We have 8n + 6 =10 To solve for n means we want n alone for that we need to get the rid of 6 from left hand side first so add -6 on both side then we get 8n = 4. now to get n alone we need to get the rid of 8 so divide both side by 8 then we get n = 4/8 n = 1/2. I hope this will help you.